OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). y = 2x + 6 y = −x + 3 y = −x + 6 y = 2x + 3

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

perpendicular slope is the reverse reciprocal of 2x. 2 = 2/1 2rev = -1/2 slope = -1/2 y-intercept = 3 y= (-1/2)x+3

OpenStudy (anonymous):

The reciprocal is just flipping a fraction and negating it.